OpenStudy (anonymous):

beth is two years older than julio. gerald is twice as old as beth. debra is twice as old as gerald. the sum of their ages is 38. how old is beth?

OpenStudy (gw2011):

Let the following letters represent the problem: B=Beth J=Julio=B-2 Julio is two years younger than Beth G=Gerald=2B D=Debra=2G=4B The equation becomes: B+(B-2)+2B+4B=38 B+B-2+2B+4B=38 8B=40 B=5 Therefore, Beth is 5 years old. From here you can find all the other ages by how the problem was set up above. Once you do this you can substitute in the numbers in the equation or in the description of the problem to prove the correctness of the answer.